OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 7, -7, 7, -7,

OpenStudy (anonymous):

@amistre64 @thomaster @UnkleRhaukus

OpenStudy (unklerhaukus):

what formula for n will make odd terms positive, and even terms negative? 1, -1, 1, -1, ...

OpenStudy (anonymous):

-1,-1

OpenStudy (unklerhaukus):

(-1)^?

OpenStudy (anonymous):

(-1)^2

OpenStudy (unklerhaukus):

it needs to be a function of n

OpenStudy (anonymous):

an = 7 • (-1)n

OpenStudy (anonymous):

(-1)^n i ment it to say

OpenStudy (unklerhaukus):

n = 1, 2 , 3, 4, ... a(n) = 7, -7, 7, -7, ...

OpenStudy (anonymous):

an = 7 • 1^n - 1

OpenStudy (unklerhaukus):

n = 1, 2, 3, ... (-1)^n = (-1)^1, (-1)^2, (-1)^3 ... = -1, 1, -1, ...

OpenStudy (anonymous):

so would it be an = 7 • (-1)^n - 1

OpenStudy (unklerhaukus):

n = 1, 2, 3, ... a_n= 7 • (-1)^(n-1) = 7 • (-1)^(1-1), 7 • (-1)^(2-1), 7 • (-1)^(3-1) .... = 7, -7 , 7 ....

OpenStudy (anonymous):

so...an = 7 • 1^n + 1

OpenStudy (anonymous):

my bad scratch that last post

OpenStudy (anonymous):

tnhx though you was very helpful